OpenStudy (anonymous):

Which of the following is/are true? I.(log[2]1) (log[4]5) = log[2]8 - log[3]27 II. log[9]2+log[9]3 > log[9]5 III. log[3](3+9) < log[6]6+log[3]4 a. II and III b. II only c. III only d. I and II

OpenStudy (***[isuru]***):

in ur third choice.... is it just log[6] ?

OpenStudy (anonymous):

sorry... I've change it its log[6] 6

OpenStudy (anonymous):

I got D

OpenStudy (***[isuru]***):

it's only 2

OpenStudy (***[isuru]***):

1st one log[2]8 = log[2]2^3 = 3log[2]2 = 3 log[3]27 = log[3]3^3 = 3log[3]3 = 3 then log[2]8 - log[3]27 = 3 - 3 = 0 but log[2]1(log[4]5) can;t be zero so it;s wrong

OpenStudy (***[isuru]***):

for the third one log[6]6 = 1 then log[6]6 + log[3]4 = 1 + log[3]4 also log[3](3 + 9) = log[3]12 = log[3](3 x4) = log[3]3 + log[3]4 = 1 + log[3]4 so... log[3](3 + 9 ) = log[6]6 + log[3]4 so the third one is also false

OpenStudy (***[isuru]***):

so.. the only answer left is 2nd one

OpenStudy (***[isuru]***):

besides log[9]2 + log[9]3 = log[9](2x3) = log[9]6 and log[9]6 > log[9]5

OpenStudy (anonymous):

thanks again !

OpenStudy (***[isuru]***):

u r welcome!